# Low-Degree Polynomials Are Good Extractors

### v1

@@
Low-Degree Polynomials Are Good Extractors
Omar Alrabiah UC Berkeley. oalrabiah@berkeley.edu. Supported by a Saudi Arabian Cultural Mission (SACM) Scholarship, NSF Award CCF-2210823, and a Simons Investigator Award (Venkatesan Guruswami)    Jesse Goodman UT Austin. jpmgoodman@utexas.edu. Supported by a Simons Investigator Award (#409864, David Zuckerman).    Jonathan Mosheiff Ben-Gurion University. mosheiff@bgu.ac.il. Supported by an Alon Fellowship.    João Ribeiro NOVA LINCS & NOVA School of Science and Technology. joao.ribeiro@fct.unl.pt. Supported in part by NOVA LINCS (ref. UIDB/04516/2020) with the financial support of FCT - Fundação para a Ciência e a Tecnologia.
@@
    Low-degree polynomials extract from small families. We show that a random low-degree polynomial is a good low-error extractor for any small family of sources. In particular, we improve the positive result of Alrabiah, Chattopadhyay, Goodman, Li, and Ribeiro (ICALP 2022) for local sources, and give new results for polynomial sources and variety sources via a single unified approach.
@@
    Low-degree polynomials extract from sumset sources. We show that a random low-degree polynomial is a good extractor for sumset sources, which are the most general large family of sources (capturing independent sources, interleaved sources, small-space sources, and more). This extractor achieves polynomially small error, and its min-entropy requirement is tight up to a square.
Our results on sumset extractors imply new complexity separations for linear ROBPs, and the tools that go into its proof have further applications, as well. The two main tools we use are a new structural result on sumset-punctured Reed-Muller codes, paired with a novel type of reduction between randomness extractors. Using the first new tool, we strengthen and generalize the extractor impossibility results of Chattopadhyay, Goodman, and Gurumukhani (ITCS 2024). Using the second, we show the existence of sumset extractors for min-entropy $k=O(\log(n/\varepsilon))$, resolving an open problem of Chattopadhyay and Liao (STOC 2022).
@@ 1 Introduction
How biased is a random degree $d$ polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$?
@@
For these definitions of random and bias, Ben-Eliezer, Hod, and Lovett [BHL12] – building on a pair of earlier papers [KS05, ABK08] – provided a complete answer to the above question. In particular, they showed sharp concentration bounds on $\operatorname{bias}(f)$, concluding that a random degree $d$ polynomial is essentially unbiased on a uniform input, with extremely high probability. More precisely, they proved the following.
@@
For every $\delta\in(0,1)$ there is a constant $c>0$ such that the following holds. Let $d\in\mathbb{N}$ be an integer satisfying $1\leq d\leq(1-\delta)k$. Then for a random degree $d$ polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$,
@@
$\operatorname{bias}_{\mathbf{X}}(f):=\Pr_{x\sim\mathbf{U}_{n}}[f(x)=0]-\Pr_{x\sim\mathbf{U}_{n}}[f(x)=1].$
Given this definition, a simple observation is that $\operatorname{bias}_{\mathbf{U}_{n}}(f)=\operatorname{bias}(f)$, and thus establishing concentration bounds for $\operatorname{bias}_{\mathbf{X}}(f)$ is a strictly more general problem than doing so for $\operatorname{bias}(f)$. But how should we use this generality? And for what distributions $\mathbf{X}$ is it actually interesting to understand $\operatorname{bias}_{\mathbf{X}}(f)$? Recall that we would like to understand $\operatorname{bias}_{\mathbf{X}}(f)$ for weakly random $\mathbf{X}$, but it is still not clear what weakly random should mean. To answer all of these questions and more, we enter the world of randomness extraction.
@@
$|\mathsf{Ext}(\mathbf{X})-\mathbf{U}_{m}|\leq\varepsilon,$
where ||$\cdot$|| ⋅ | denotes statistical distance.
@@
$\operatorname{bias}_{\mathbf{X}}(\mathsf{Ext})\leq 2\varepsilon.$
Thus, returning to our discussion at the beginning of the paper, we see that getting good bounds on $\operatorname{bias}_{\mathbf{X}}(f)$ for a random low-degree polynomial $f$ is equivalent to showing that a random low-degree polynomial is a good extractor for $\mathbf{X}$. Thus, in order to figure out interesting distributions $\mathbf{X}$ for which to pursue upper bounds on $\operatorname{bias}_{\mathbf{X}}(f)$, it is only natural to borrow some motivation from extractor theory. We do so, below.
@@
In order to show that a random low-degree polynomial extracts from a distribution $\mathbf{X}\sim\{0,1\}^{n}$ (equivalently, in order to upper bound $\operatorname{bias}_{\mathbf{X}}(f)$), it is easy to see that an absolute minimum requirement is that $\mathbf{X}$ contains some “randomness.” To formalize this notion, it is standard to use min-entropy, defined as follows.
@@
If we can show that $\operatorname{bias}_{\mathbf{X}}(f)$ is low with extremely high probability over $f$, then we can conclude that random low-degree polynomials extract from any small family $\mathcal{X}$ of sources. Importantly, many well-studied families of sources are, in fact, very small. In particular, this is true of all families for which random low-degree polynomial extractors have been studied: uniform sources [BHL12], affine sources [CT15], and local sources [ACG+22]. [^3] The family of uniform sources is the trivial family $\mathcal{X}=\{\mathbf{U}_{n}\}$, while the faimly of affine sources (of min-entropy $k$) consists of all $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ that are uniform over a $k$-dimensional affine subspace of $\mathbb{F}_{2}^{n}$. Local sources, on the other hand, consist of sources of the form $\mathbf{X}=f(\mathbf{U}_{\ell})$, where $f$ is some function where each output bit depends on a bounded number of input bits. Thus, showing that a random low-degree polynomial is unbiased on any single source $\mathbf{X}$ of min-entropy $k$ could lead to a result that subsumes (and greatly generalizes) all previous work.
@@
In the remainder of this paper, our goal is to answer both of the questions presented above. In doing so, we hope to provide new insight into the power of a fundamental computational model (low-degree polynomials) for a fundamental computational task (extracting randomness). As it turns out, however, there are a few other reasons why answering these questions might be useful. Before we formally present our main results, we highlight some of these, below.
@@
Low-complexity extractors have important applications in the real world and theory. In the real world, low-complexity extractors are more likely to be implemented and exhibit a reasonable running time. In theory, low-complexity extractors serve as fundamental building blocks in the construction of key cryptographic [Lu02, Vad04] and pseudorandom primitives [Li11, LZ23]. [^4] In fact, several well-known explicit extractors are themselves low-degree polynomials [CG88, BSHR+01, Bou05]! Their study has also led to important structural results for well-studied families of distributions [CT15, ACG+22]. Because of this, low-complexity extractors have received a lot of attention in the literature [DM02, Lu02, Vad04, Vio05, DT09, Li11, BG13, GVW15, CT15, CL18, ACG+22, HIV22, CW24, LZ23], with the works of Cohen and Tal [CT15] and Alrabiah, Chattopadhyay, Goodman, Li, and Ribeiro [ACG+22] specifically focusing on the power of random low-degree polynomials as extractors (for affine sources and local sources, respectively).
@@
Finally, low-complexity extractors can help us establish fine-grained complexity separations (as advocated for in, e.g., [HHTT23]). In more detail, extractors are known to exhibit strong lower bounds against a variety of well-studied complexity classes, including low-depth circuits [HR15, GKW21], general circuits [DK11, GK16, FGHK16, GKW21, LY22], various flavors of branching programs [HR15, CGZ22, GPT22, CL23], and more [CS16, GKW21]. Showing that there exist extractors in a low-level complexity class $\mathcal{C}$ would allow one to separate $\mathcal{C}$ from the classes above.
@@ 1.1.1 Low-degree polynomials extract from small families
In order to prove that random low-degree polynomials can extract from any small family, we first show that a random low-degree polynomial extracts from a single source. In particular, we prove the following, which can be viewed as (1) a “low-degree” version of the classical fact that a random function extracts from a single source [Vad12, Proposition 6.12], and (2) a generalization of the result that a random low-degree polynomial has low bias [BHL12, Lemma 1.2] (Theorem ).
@@
$\Pr_{f}\left[\operatorname{bias}_{\mathbf{X}}(f)>2^{-ck/d}\right]\leq 2^{-c\binom{k}{\leq d}}.$
@@
We also note that by a standard application of the XOR lemma [DLWZ14, Lemma 3.8], it is straightforward to extend Theorem 1 to show that a sequence of independent, uniformly random degree $d$ polynomials $f_{1},f_{2},\dots,f_{m}\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ can be concatenated to create a multi-bit extractor for $\mathbf{X}$ [^7] In order to apply the XOR lemma, the only observation needed is that the XOR of (any number of) independent, uniformly random degree $d$ polynomials applied to $\mathbf{X}$ is, itself, a uniformly random degree $d$ polynomial applied to $\mathbf{X}$. In fact, this can further be extended to show that the sequence $f_{1},f_{2},\dots,f_{m}$ not only extracts $m$ uniform bits from $\mathbf{X}$, but has low correlation with any (short) fixed function $g$ applied to $\mathbf{X}$ [^8] This can be done by first conditioning on the output of $g(\mathbf{X})$, which will only cause $\mathbf{X}$ to lose a little bit of min-entropy. Finally, if we set $k=n$, $m=1$, and $g$ to have output length $1$, this result can be interpreted as nontrivial bounds on the list-size of Reed-Muller codes at the extreme (relative) radius of $1/2-2^{-\Omega(n/d)}$. This appears to be the first result of this form (c.f. [KLP12, ASW15, ASY20]), and naturally extends to punctured Reed-Muller codes (by picking $k<n$).
@@
For every $\delta>0$ there exists a constant $c>0$ such that the following holds for all positive integers $n\geq k$ and $d\leq(1-\delta)k$. For any family $\mathcal{X}$ of $(n,k)$-sources with size $|\mathcal{X}|<2^{c\binom{k}{\leq d}}$, a random degree $d$ polynomial $f\colon\{0,1\}^{n}\to\{0,1\}$ is a $2^{-ck/d}$-extractor for $\mathcal{X}$, except with probability at most $2^{-c\binom{k}{\leq d}}$.
@@
  • Local sources: There exists a degree $\leq d$ polynomial $f\colon\{0,1\}^{n}\to\{0,1\}$ that is a $2^{-ck/d}$-extractor for $r$-local sources with min-entropy
@@
  • Polynomial sources: There exists a degree $\leq d$ polynomial $f\colon\{0,1\}^{n}\to\{0,1\}$ that is a $2^{-ck/d}$-extractor for degree $r$ polynomial sources with min-entropy
@@
  • Variety sources: There exists a degree $\leq d$ polynomial $f\colon\{0,1\}^{n}\to\{0,1\}$ that is a $2^{-ck/d}$-extractor for degree $r$ variety sources with min-entropy
@@
We make a few brief remarks about this result. First, we note that our result on local sources significantly improves the parameters of the previous best result [ACG+22, Theorem 1.1], which required min-entropy $k\geq C2^{r}r^{2}d(2^{r}n\log n)^{1/d}$ and had error $\varepsilon=2^{-ck/(d^{3}2^{r}r^{2})}$. Second, we highlight that our result on polynomial sources is rather surprising, as it shows that polynomials can be used to extract from polynomial sources. Perhaps this can be used to make progress on (the extremely challenging goal of) constructing explicit extractors for polynomial sources [DGW09, CGG24], as it suggests that it is possible to improve the quality of a polynomial source while keeping it as a polynomial source. Third, we mention one interesting application of our result on variety sources. By exploiting known hardness results for variety extractors [LY22, GK16, GKW21], it follows that (general) circuits of size $3.1n-o(n)$ cannot compute (all) degree $3$ polynomials, circuits of size $3.11n-o(n)$ cannot compute degree $4$ polynomials, and circuits of size $3.9n-o(n)$ cannot compute degree $18$ polynomials. Finally, we note that in Section 4.3, we show that Corollary 1 can be used to prove that random low-degree polynomials are good linear seeded extractors. [^9] Recall that a linear seeded extractor only needs to be linear on the source (fixing the seed), but may be an arbitrarily high-degree polynomial in general.
@@ 1.1.2 Low-degree polynomials extract from sumset sources
In the second main part of our paper, we show that random low-degree polynomials are also good extractors for sumset sources, which are the most general well-studied large family of sources. We prove the following, and later show that this result can be extended to obtain near-optimal existential results for sumset extractors.
@@
For any $\delta\in(0,1)$ there exist absolute constants $C,c>0$ such that for any positive integers $n$, $d\leq\frac{c\log n}{\log\log n}$, min-entropy $k\geq Cdn^{\frac{2(1+\delta)}{d}}$, and error $\varepsilon=n^{-c}$, a random degree $d$ polynomial $f\in\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ is a $\varepsilon$-extractor for $(n,k)$-sumset sources, with probability at least $1-2^{-c\binom{k/6}{d/2}}$. For the easier setting of dispersers, [^10] A disperser is an extractor $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ with nontrivial error $\varepsilon<1/2$. we can improve the min-entropy requirement to $k\geq Cdn^{2/d}$, and have this hold with probability at least $1-2^{-c\binom{k}{\leq d}}$.
A few remarks are in order. First, while our results for small families hold for exponentially small error, our result for sumset sources only holds for polynomially small error. On the other hand, similar to Theorem 1, Theorem 2 is (nearly) tight, as Cohen and Tal [CT15] show that degree $d$ polynomials cannot extract from affine sources (and thus sumset sources) with min-entropy below $k=cdn^{\frac{1}{d-1}}$. Finally, we remark that our sumset extractors have interesting consequences for linear read-once branching programs, and our proof requires two key new ingredients, each of which has a “bonus” application in pseudorandomness.
In more detail, linear read-once branching programs (ROBPs) are a new type of computational model [GPT22], which simultaneously generalize both standard ROBPs and parity decision trees. At each point in the branching program, instead of querying a single input variable, the ROBP is allowed to query an arbitrary linear function of the input (so long as it is linearly independent of all previous queries). We observe that by leveraging standard results on finite fields [LN97, Lemma 6.21] (see also [BV10, Lemma 17]), linear ROBPs of constant width $w=O(1)$ can compute any polynomial of degree $2$. On the other hand, Theorem 2 (combined [CL23, Theorem 1]) implies that linear ROBPs require exponential width $w=2^{n-o(n)}$ to compute polynomials of degree $3$. This is a huge, perhaps surprising, jump in complexity. [^11] Indeed, one might expect the width $w$ to somehow grow proportionately with the degree $d$ of the polynomial that must computed. However, this shows that the width jumps from constant to exponential, simply by moving from degree $2$ to degree $3$.
Finally, we highlight that the proof of Theorem 2 requires two new ingredients - each of which find a “bonus” application in pseudorandomness, and may therefore be of independent interest. The first key ingredient is a new result about the structure of Reed-Muller codes that are punctured on sumsets. As a bonus application, we use this to build new “evasive sets,” which we then use to improve the extractor impossibility results of Chattopadhyay, Goodman, and Gurumukhani [CGG24], and to get a low-error version of Theorem 2 for the special case of degree $4$ polynomials and independent sources. Interestingly, the latter result relies on the recent breakthrough resolution of the PFR conjecture from additive combinatorics [GGMT23].
The second key ingredient in the proof of Theorem 2 is a novel type of reduction between extractors. While most reductions between extractors rely on showing that a source can be equipped with structure by breaking it down into a convex combination of well-behaved distributions via a deterministic process, we show that doing so via a randomized process can sometimes make this task much easier. As a bonus application, we establish the first near-optimal existential result for sumset extractors.
Theorem 3 (Existential result for sumset extraction).
@@
For every constant $\gamma>0$ there exists a constant $C>0$ such that the following holds for any integer $n$ and $\varepsilon>0$: With probability at least $0.99$, a uniformly random function $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is a $(k,\varepsilon)$-sumset extractor for $k=(2+\gamma)(\log n+2\log(1/\varepsilon))+C$.
In particular, we show that there exist sumset extractors for min-entropy $k=O(\log(n/\varepsilon))$, resolving an open problem of Chattopadhyay and Liao [CL22].
@@
The remainder of the paper is organized as follows. In Section 2, we provide an overview of our main techniques. Then, in Section 3, we record some basic preliminaries on notation, probability, and extractors. The first main technical part of this paper is Section 4, where we show that random low-degree polynomials extract from small families, and prove all results listed in Section 1.1.1. The second main technical part of the paper consists of Sections 5 and 6, where we prove all results listed in Section 1.1.2. In particular, we show that random low-degree polynomials extract from sumset sources in Section 5, where we also provide our general existential results for sumset extractors. In Section 6, we show that one the key ingredients from the prior section can be used to establish the existence of “evasive sets,” which are then used to obtain a few additional results. Finally, we wrap up with some open problems in Section 7.

### v2

@@
Abstract
@@
    Low-degree polynomials extract from small families. We show that a random low-degree polynomial is a good low-error extractor for any small family of sources. In particular, we improve the positive result of Alrabiah, Chattopadhyay, Goodman, Li, and Ribeiro (ICALP 2022) for local sources, and give new results for polynomial and variety sources via a single unified approach.
@@
    Low-degree polynomials extract from sumset sources. We show that a random low-degree polynomial is a good extractor for sumset sources, which are the most general large family of sources (capturing independent sources, interleaved sources, small-space sources, and more). Formally, for any even $d$, we show that a random degree $d$ polynomial is an $\varepsilon$-error extractor for $n$-bit sumset sources with min-entropy $k=O(d(n/\varepsilon^{2})^{2/d})$. This is nearly tight in the polynomial error regime.
Our results on sumset extractors imply new complexity separations for linear ROBPs, and the tools that go into its proof may be of independent interest. The two main tools we use are a new structural result on sumset-punctured Reed-Muller codes, paired with a novel type of reduction between extractors. Using the new structural result, we obtain new limits on the power of sumset extractors, strengthening and generalizing the impossibility results of Chattopadhyay, Goodman, and Gurumukhani (ITCS 2024).
@@ 1 Introduction
How biased is a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$?
@@
For these definitions of random and bias, Ben-Eliezer, Hod, and Lovett [BHL12] – building on a pair of earlier papers [KS05, ABK08] – provided a complete answer to the above question. In particular, they showed sharp concentration bounds on $|\operatorname{bias}(f)|$, concluding that a random degree $d$ polynomial is essentially unbiased on a uniform input, with extremely high probability. More precisely, they proved the following.
@@
For every $\delta\in(0,1)$ there is a constant $c>0$ such that the following holds. Let $d\in\mathbb{N}$ be an integer satisfying $1\leq d\leq(1-\delta)k$. Then for a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$,
@@
$\operatorname{bias}_{\mathbf{X}}(f):=\Pr_{x\sim\mathbf{X}}[f(x)=0]-\Pr_{x\sim\mathbf{X}}[f(x)=1].$
Given this definition, a simple observation is that $\operatorname{bias}_{\mathbf{U}_{n}}(f)=\operatorname{bias}(f)$, and thus establishing concentration bounds for $|\operatorname{bias}_{\mathbf{X}}(f)|$ is a strictly more general problem than doing so for $|\operatorname{bias}(f)|$. But how should we use this generality? And for what distributions $\mathbf{X}$ is it actually interesting to understand $|\operatorname{bias}_{\mathbf{X}}(f)|$? Recall that we would like to understand $|\operatorname{bias}_{\mathbf{X}}(f)|$ for weakly random $\mathbf{X}$, but it is still not clear what weakly random should mean. To answer all of these questions and more, we enter the world of randomness extraction.
@@
$\Delta(\mathsf{Ext}(\mathbf{X}),\mathbf{U}_{m})\leq\varepsilon,$
where $\Delta(\cdot,\cdot)$ denotes statistical distance. For short, we also call $\mathsf{Ext}$ an $\varepsilon$-extractor for $\mathcal{X}$.
@@
$|\operatorname{bias}_{\mathbf{X}}(\mathsf{Ext})|\leq 2\varepsilon.$
Thus, returning to our original discussion, we see that getting good bounds on $|\operatorname{bias}_{\mathbf{X}}(f)|$ for a random low-degree polynomial $f$ is equivalent to showing that a random low-degree polynomial is a good extractor for $\mathbf{X}$. Thus, in order to figure out interesting distributions $\mathbf{X}$ for which to pursue upper bounds on $|\operatorname{bias}_{\mathbf{X}}(f)|$, it is only natural to borrow some motivation from extractor theory. We do so, below.
@@
In order to show that a random low-degree polynomial extracts from a distribution $\mathbf{X}\sim\{0,1\}^{n}$ (equivalently, in order to upper bound $|\operatorname{bias}_{\mathbf{X}}(f)|$), it is easy to see that an absolute minimum requirement is that $\mathbf{X}$ contains some “randomness.” To formalize this notion, it is standard to use min-entropy, defined as follows.
@@
If we can show that $|\operatorname{bias}_{\mathbf{X}}(f)|$ is low with extremely high probability over $f$, then we can conclude that random low-degree polynomials extract from any small family $\mathcal{X}$ of sources. Importantly, many well-studied families of sources are, in fact, very small. In particular, this is true of all families for which random low-degree polynomial extractors have been studied: uniform sources [BHL12], affine sources [CT15], and local sources [ACG+22]. [^3] The family of uniform sources is the trivial family $\mathcal{X}=\{\mathbf{U}_{n}\}$, while the family of affine sources (of min-entropy $k$) consists of all $\mathbf{X}\sim\mathbb{F}_{2}^{n}$ that are uniform over a $k$-dimensional affine subspace of $\mathbb{F}_{2}^{n}$. Local sources, on the other hand, consist of sources of the form $\mathbf{X}=f(\mathbf{U}_{\ell})$, where $f$ is some function where each output bit depends on a bounded number of input bits. Thus, showing that a random low-degree polynomial is unbiased on any single source $\mathbf{X}$ of min-entropy $k$ could lead to a result that subsumes (and greatly generalizes) all previous work.
@@
In the remainder of this paper, our goal is to answer both of the questions presented above. In doing so, we hope to provide new insight into the power of a fundamental computational model (low-degree polynomials) for a fundamental computational task (randomness extraction). As it turns out, however, there are a few other reasons why answering these questions might be useful. Before we formally present our main results, we highlight some of these, below.
@@
[…]
@@
Finally, low-complexity extractors can help us establish fine-grained complexity separations (as advocated for in, e.g., [HHTT23]). In more detail, extractors are known to exhibit strong lower bounds against a variety of well-studied complexity classes, including low-depth circuits [HR15, GKW21], general circuits [DK11, GK16, FGHK16, GKW21, LY22], various flavors of branching programs [HR15, CGZ22, GPT22, CL23], and more [CS16, GKW21, GG25, CG25]. Showing that there exist extractors in a low-level complexity class $\mathcal{C}$ would allow one to separate $\mathcal{C}$ from the classes above.
@@ 1.1.1 Low-degree polynomials extract from small families
In order to prove that random low-degree polynomials can extract from any small family, we first show that a random low-degree polynomial extracts from a single source. In particular, we prove the following, which can be viewed as both (1) a “low-degree” version of the classical fact that a random function extracts from a single source [Vad12, Proposition 6.12], and (2) a generalization of the result that a random low-degree polynomial has low bias [BHL12, Lemma 1.2] (Theorem ).
@@
$\Pr_{f}\left[|\operatorname{bias}_{\mathbf{X}}(f)|>2^{-ck/d}\right]\leq 2^{-c\binom{k}{\leq d}}.$
@@
We also note that by a standard application of the XOR lemma [DLWZ14, Lemma 3.8], it is straightforward to extend Theorem 1 to show that a sequence of independent, uniformly random degree $d$ polynomials $f_{1},f_{2},\dots,f_{m}:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ can be concatenated to create a multi-bit extractor for $\mathbf{X}$ [^7] In order to apply the XOR lemma, the only observation needed is that the XOR of (any number of) independent, uniformly random degree $d$ polynomials applied to $\mathbf{X}$ is, itself, a uniformly random degree $d$ polynomial applied to $\mathbf{X}$. In fact, this can further be extended to show that the sequence $f_{1},f_{2},\dots,f_{m}$ not only extracts $m$ uniform bits from $\mathbf{X}$, but has low correlation with any (short) fixed function $g$ applied to $\mathbf{X}$ [^8] This can be done by first conditioning on the output of $g(\mathbf{X})$, which will only cause $\mathbf{X}$ to lose a little bit of min-entropy. Finally, if we set $k=n$, $m=1$, and $g$ to have output length $1$, this result can be interpreted as nontrivial bounds on the list-size of Reed-Muller codes at the extreme (relative) radius of $1/2-2^{-\Omega(n/d)}$. This appears to be the first result of this form (c.f. [KLP12, ASW15, ASY20]), and naturally extends to punctured Reed-Muller codes (by picking $k<n$).
@@
For every $\delta>0$ there exists a constant $c>0$ such that for all $n\geq k\geq d\in\mathbb{N}$ with $d\leq(1-\delta)k$, the following holds. For any family $\mathcal{X}$ of $(n,k)$-sources with size $|\mathcal{X}|<2^{c\binom{k}{\leq d}}$, a random degree $d$ polynomial $f:\{0,1\}^{n}\to\{0,1\}$ is a $2^{-ck/d}$-extractor for $\mathcal{X}$, except with probability at most $2^{-c\binom{k}{\leq d}}$.
@@
  • Local sources: There exists a degree $\leq d$ polynomial $f:\{0,1\}^{n}\to\{0,1\}$ that is a $2^{-ck/d}$-extractor for $r$-local sources with min-entropy
@@
  • Polynomial sources: There exists a degree $\leq d$ polynomial $f:\{0,1\}^{n}\to\{0,1\}$ that is a $2^{-ck/d}$-extractor for degree $r$ polynomial sources with min-entropy
@@
  • Variety sources: There exists a degree $\leq d$ polynomial $f:\{0,1\}^{n}\to\{0,1\}$ that is a $2^{-ck/d}$-extractor for degree $r$ variety sources with min-entropy
@@
We make a few brief remarks about this result. First, we note that our result on local sources significantly improves the parameters of the previous best result [ACG+22, Theorem 1.1], which required min-entropy $k\geq C2^{r}r^{2}d(2^{r}n\log n)^{1/d}$ and had error $\varepsilon=2^{-ck/(d^{3}2^{r}r^{2})}$. Second, we highlight that our result on polynomial sources may be surprising, as it shows that polynomials can be used to extract from polynomial sources. Perhaps this can be used to make progress on (the challenging goal of) constructing explicit extractors for polynomial sources [DGW09, CGG24], as it suggests that it is possible to improve the quality of a polynomial source while keeping it as a polynomial source. Third, we mention one interesting application of our result on variety sources. By exploiting known hardness results for variety extractors [LY22, GK16, GKW21], it follows that (general) circuits of size $3.1n-o(n)$ cannot compute (all) degree $3$ polynomials, circuits of size $3.11n-o(n)$ cannot compute degree $4$ polynomials, and circuits of size $3.9n-o(n)$ cannot compute degree $18$ polynomials. Finally, we note that in Section 4.3, we show that Corollary 1 (in fact, Theorem ) can be used to prove low-degree polynomials are good linear seeded extractors. [^9] Recall that a linear seeded extractor only needs to be linear on the source (fixing the seed), but may be an arbitrarily high-degree polynomial in general.
@@ 1.1.2 Low-degree polynomials extract from sumset sources
In the second main part of our paper, we show that random low-degree polynomials are also good extractors for sumset sources, which are the most general well-studied large family of sources.
@@
There exists a constant $C>0$ such that for any $n\geq k\geq d\in\mathbb{N}$ and $\varepsilon>0$ such that $k\geq Cd(n/\varepsilon^{2})^{1/\lfloor d/2\rfloor}$, a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is an $\varepsilon$-extractor for $(n,k)$-sumset sources, with probability at least $1-2^{-\varepsilon^{2}\binom{k/C}{2\lfloor d/2\rfloor}}\geq 1-2^{-n^{2}/\varepsilon^{2}}$.
We highlight a few key specializations of the above theorem – focusing on even $d$, for simplicity. First, we note that in the disperser regime, [^10] A disperser is an extractor $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ with nontrivial error $\varepsilon<1/2$. Theorem 2 shows that there exist degree $\leq d$ polynomials that disperse from sumset sources with min-entropy $k=O(dn^{2/d})$. This is nearly tight, since Cohen and Tal [CT15] show that degree $\leq d$ polynomials cannot extract from affine sources (and thus sumset sources) with min-entropy below $k=\Omega(dn^{1/(d-1)})$. (In Section 5.3, we show the same impossibility result holds for independent sources, which are the other special case of sumset sources.) Second, we note that in the polynomial error regime $\varepsilon=n^{-\gamma}$, Theorem 2 shows that there exist degree $\leq d$ polynomials that are $\varepsilon$-extractors for sumset sources with min-entropy $k=O(dn^{2(1+2\gamma)/d})$. Third, in the arbitrary error regime $\varepsilon>0$, Theorem 2 shows that a random degree $d=O(\log(n/\varepsilon))$ polynomial $f$ is an $\varepsilon$-extractor for sumset sources with min-entropy $k=O(\log(n/\varepsilon))$. This strengthens the existential result of Mrazović [Mra16], who obtained such a min-entropy requirement for a uniformly random function $f$.
Finally, we highlight that our sumset extractors have interesting consequences for linear read-once branching programs, and our proof requires two new tools, which may be of independent interest.
Linear ROBPs
In more detail, linear read-once branching programs (ROBPs) are a new type of computational model [GPT22], which simultaneously generalize both standard ROBPs and parity decision trees. At each point in the branching program, instead of querying a single input variable, the ROBP is allowed to query an arbitrary linear function of the input (so long as it is linearly independent of all previous queries). We observe that by leveraging standard results on finite fields [LN97, Lemma 6.21] (see also [BV10, Lemma 17]), linear ROBPs of constant width $w=O(1)$ can compute any polynomial of degree $2$. [^11] In more detail, [BV10, Lemma 17] asserts that for any quadratic $q:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$, there is some $B\in\mathbb{F}_{2}^{m\times n}$ with full row rank and some affine $L:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ such that $q(x)=\langle(Bx)_{\leq m/2},(Bx)_{>m/2}\rangle+L(x)$. Since each row in $B$ is linearly independent, the inner product can be computed using $\leq 2$ bits of storage. Then, using $\leq 1$ additional bit of storage (in case $L$ is linearly dependent on the rows in $B$), one can simultaneously compute $L(x)$. Thus, $q(x)$ can be computed by a constant-width linear ROBP. On the other hand, Theorem 2 (combined with [CL23, Theorem 1]) implies that linear ROBPs require exponential width $w=2^{n-o(n)}$ to compute polynomials of degree $4$. This is a huge, perhaps surprising, jump in complexity. [^12] Indeed, one might expect the width $w$ to somehow grow proportionately with the degree $d$ of the polynomial that must computed. However, this shows that the width jumps from constant to exponential, simply by moving from degree $2$ to degree $4$.
Sumset-punctured Reed-Muller codes
Finally, we highlight that the proof of Theorem 2 requires two new key ingredients, which may be of independent interest. The first key ingredient is a new result about the structure of Reed-Muller codes that are punctured on sumsets. As a bonus application, we use this to build new “evasive sets,” which we then use to improve the extractor impossibility results of Chattopadhyay, Goodman, and Gurumukhani [CGG24], and to get a low-error version of Theorem 2 for the special case of degree $4$ polynomials and independent sources. Interestingly, the latter result relies on the recent breakthrough resolution of Marton’s PFR conjecture from additive combinatorics [GGMT23].
A novel reduction between extractors
The second key ingredient in the proof of Theorem 2 is a novel type of reduction between extractors. While most reductions between extractors rely on showing that a source can be equipped with structure by breaking it down into a convex combination of well-behaved distributions via a deterministic process, we show that doing so via a careful (correlated) randomized process can sometimes make this task much easier. In Appendix A, we illustrate a simpler variant of this idea in order to give an alternative proof that a uniformly random function is an extractor for sumset sources with min-entropy $k=O(\log(n/\varepsilon))$ – a result first established by Mrazović [Mra16].
@@
The remainder of the paper is organized as follows. In Section 2, we provide an overview of our main techniques. Then, in Section 3, we record some basic preliminaries on notation, probability, and extractors. The first main technical part of this paper is Section 4, where we show that random low-degree polynomials extract from small families, and prove all results listed in Section 1.1.1. The second main technical part of the paper consists of Section 5, where we prove our main result from Section 1.1.2. In Section 6, we show that one of the key ingredients from the prior section can be used to establish the existence of “evasive sets,” which are then used to obtain a few additional results. (In Appendix A, we show that the other key ingredient can be used to obtain an alternative proof of the result by Mrazović [Mra16].) Finally, we wrap up with some open problems in Section 7.